OpenStudy (anonymous):

which of the following statements is (are) true for all the positive integers? 1) n^3+2n is divisible by 3 2) 5^(2n)-1 is divisible by 24

OpenStudy (anonymous):

finally someone looks at it

mathslover (mathslover):

I think both are correct dude

mathslover (mathslover):

I just verified them with the help of a calculator

OpenStudy (anonymous):

how do you do it on the calculatyor?

OpenStudy (anonymous):

i have the ti-84 plus

mathslover (mathslover):

for 1) n^3+2n is divisible by 3 let n = 3 3^3 + 2(3) = 27 + 6 = 33 that is divisible by 3 now let us take a big number n = 24 3^24 +2(24) is divisble by 3

OpenStudy (anonymous):

yeah i did that i got 33 then what?

OpenStudy (anonymous):

now let us take a big number n = 24 3^24 +2(24) is divisble by 3 i dont get this?

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhh gotcha

OpenStudy (anonymous):

swagggg thanks man

mathslover (mathslover):

ok.. best of luck

OpenStudy (barrycarter):

Hint: ((n+1)^3+2(n+1)) - (n^3+2n) = 3*(1+n+n^2) In other words, n^3+2n increases by a multiple of 3 each time you increase n. So, if it's true for n=1 (which it is), it's true for all n. Similarly, (5^(2(n+1))-1)-(5^(2n)-1) simplifies to 24*25^n, so, when you increase n, you increase 5^(2n)-1 by a multiple of 24.